DISSOCIATION DYNAMICS OF HYDROGEN-DIHALOGEN COMPLEXES

\underline{R. A. LOOMIS}, Department of Chemistry, Washington University, Saint Louis, Missouri, 63130.

Two-color, pump-probe spectroscopy has been used to investigate the vibrational predissociation dynamics of numerous conformers of hydrogen-dihalogen complexes. Specifically, different H_{2}-ICl(B,v), D_{2}-ICl(B,v), H_{2}-I_{2}(B,v) and D_{2}-I_{2}(B,v) intermolecular vibrational levels are prepared. The subsequent ICl(B,v) and I_{2}(B,v) product-state distributions are then measured with rotational-state resolution. The roles of the geometry of the initially prepared complex and of ortho- and para-H_{2} on the dynamics are extracted. Similar experiments were also performed to measure the binding energies of all of the conformers.